OpenStudy (anonymous):

Solve the inequality. Then graph your solution. x-1 greater than or equal to -9

OpenStudy (anonymous):

\[x-1\ge-9\]

OpenStudy (anonymous):

\[x \ge-8\]

OpenStudy (anonymous):

no sorry less than or equal to

OpenStudy (anonymous):

understand how I did that?

OpenStudy (anonymous):

\[x \le-8\]

OpenStudy (anonymous):

\[x \le -9+1\]

OpenStudy (anonymous):

I added 1 to each side, so I could isolate (solve) for x.